OpenStudy (anonymous):

Jeff’s par scores for five rounds of a golf game were 4, –3, 5, –3, and –4. What was his final score? A. –1 B. 7 C. 19 D. –19

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

I don't know how to do this

OpenStudy (anonymous):

Is it -1

OpenStudy (anonymous):

I think I'm right

OpenStudy (anonymous):

It's A.

OpenStudy (anonymous):

4 + (-3) + 5 + (-3) + (-4) = 4 - 3 + 5 - 3 - 4 = 9 - 10 = -1

OpenStudy (anonymous):

ya I'm right

OpenStudy (anonymous):

good job :)